a university law school accepts 3 out of 10 applicants if the school accepted 270 students find how many applications did they recived
Answer by jim_thompson5910(35256)

x = number total who applied



(number accepted)/(number total) = (number accepted)/(number total)

3/10 = 270/x

3*x = 10*270

3x = 2700

x = 2700/3

x = 900

So there were a total of 900 applications.